### v2

# Exploring Kolmogorov-Arnold networks for realistic image sharpness assessment

Abstract
Score prediction is crucial in realistic image sharpness assessment after informative features are collected. Recently, Kolmogorov-Arnold networks (KANs) have been developed and witnessed remarkable success in data fitting. This study presents Taylor series based KAN (TaylorKAN). Then, different KANs are explored on four realistic image databases (BID2011, CID2013, CLIVE, and KonIQ-10k) for score prediction by using 15 mid-level features and 2048 high-level features. When setting support vector regression as the baseline, experimental results indicate KANs are generally better or competitive, TaylorKAN is the best on three databases using mid-level feature input, while KANs are inferior on CLIVE when high-level features are used. This is the first study that explores KANs for image quality assessment. It sheds lights on how to select and improve KANs on related tasks.
Index Terms:
Kolmogorov-Arnold network, image sharpness assessment, image quality, machine learning.
I Introduction
Blind image sharpness/blurriness assessment (BISA) plays an important role in media quality assurance. It could automate real-time processing of large volumes of images without any reference information, allow for dynamic adjustment of image sharpness and visual fidelity in video streaming, and enhance the quality of user experience [1]. Specifically, BISA can guide the image restoration process by offering a metric to optimize algorithms for sharpening and deblurring [2].
Score prediction is indispensable when informative features have been prepared for image sharpness presentation. Support vector regression (SVR) and multi-layer perceptron (MLP) are preferred. Li $et$ $al$ craft multi-scale sharpness-aware features, and SVR performs score rating [3]. Yu $et$ $al$ design a convolutional neural network (CNN) for blurriness estimation [4], and SVR and MLP are compared [5]. Tang $et$ $al$ quantify image sharpness from intrinsic structures, and random forest is used for feature pooling and quality rating [6]. Liu $et$ $al$ craft many statistical features, and a three-hidden-layered MLP scores image quality [7]. Liu $et$ $al$ design orientation-aware features for SVR-based score prediction [8]. Chen $et$ $al$ weight local binary pattern features in spatial domain and entropy and gradient features in spectral domain, and SVR predicts the quality scores from perception features [9]. Yu $et$ $al$ construct mid-level features, and MLP, SVR, general regression neural network and elastic net are evaluated [10].
For score prediction, a CNN can be treated as an image-based feature extractor followed by a MLP in an optimization procedure. Many CNNs have been proposed. Zhu $et$ $al$ retrieve prior knowledge shared among distortions, and the network is then fine-tuned for quality prediction [11]. Song $et$ $al$ integrate domain knowledge of human visual system and natural scene statistics into a deep network for scoring [12]. Li $et$ $al$ utilize a pre-trained network to quantify the discrepancy and image quality [13]. Huang $et$ $al$ investigate the inherent relationship between the attributes and the category via graph convolution network for semantic attribute reasoning and image quality estimation [14]. Zhang $et$ $al$ optimize a deep CNN on multiple databases by a hinge constraint on learning uncertainty [15]. Li and Huo consider global and local features in multi-scale distortion perception and introduce the feedback mechanism [16]. Chen $et$ $al$ design multi-scale spatial pyramid pooling that combines block attention and image understanding for improved accuracy and generalization [17]. Sun $et$ $al$ extract low-level visual features and high-level semantic information and a staircase structure is designed for hierarchical feature integration and quality-aware representation [18]. Zhao $et$ $al$ enable learning representation by a pretext task in a self-supervised learning manner, and a quality-aware contrastive loss is employed to learn distortions from image patches [19]. Zhang $et$ $al$ develop a multitask learning scheme for image quality assessment, image scene classification and distortion type identification, and the model parameter sharing and the loss weighting are determined automatically [20]. Wu $et$ $al$ fuse multi-stage semantic features for image quality prediction, and the features are rectified using a novel multi-level channel attention module [21].
Inspired by the Kolmogorov-Arnold theorem (KAT), a novel module called Kolmogorov-Arnold Networks (KAN) has been proposed [22]. Except for the activation functions as MLP on nodes, KAN appends learnable activation functions on edges between the nodes of successive layers. Improved representation capacity has been witnessed in function approximation, data fitting, knowledge representation and collaborative filtering [23, 24, 25]. Later, different KAN variants are implemented using different mathematical functions as the activation functions on edges, such as FourierKAN-GCF [25], BSRBF_KAN [26], ChebyKAN [27], and WavKAN [28].
Despite remarkable success in knowledge representation and data fitting, little is known about KANs for score prediction in BISA. To bridge this gap, this study explores different KANs for image blurriness estimation. Firstly, Taylor series based KAN (TaylorKAN) is proposed. Then, mid-level features and deeply learned features of four realistic databases are prepared. After that, MLP, SVR and six KANs are evaluated. Results indicate that KANs are generally better than SVR and MLP, and TaylerKAN is the best on three databases when mid-level features are used.
II KAT-inspired networks
KAT asserts that a continuous multivariate function can be represented as a finite sum of continuous uni-variate functions [29]. Assuming $f:[0,1]^{n}\to\mathbb{R}$ be a continuous function, there exist continuous uni-variate functions $\phi_{q,p}$ and $\Phi_{q}$ that,
$f(x_{1},x_{2},\ldots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$ (1)
II-A The first KAN model
KAT-inspired KAN treats a multivariate function as learnable uni-variate spline functions on edges [22]. A KAN layer can be defined by a matrix of uni-variate functions $\Phi=\{\phi_{q,p}\}$ in which $p=1,2,\ldots,n_{\text{in}}$ and $q=1,2,\ldots,n_{\text{out}}$. Here, $n_{\text{in}}$ is the input dimension, $n_{\text{out}}$ is the output dimension, and $\phi_{q,p}$ is a learnable function. The activation of each node in layer $l+1$ can be computed as
$x_{l+1,j}=\sum_{i=1}^{n_{l}}\phi_{l,j,i}(x_{l,i}).$ (2)
And thus, the $L$-layered KAN can be generally described as
$\text{KAN}(x)=(\Phi_{L-1}\circ\Phi_{L-2}\circ\ldots\circ\Phi_{0})(x),$ (3)
and implemented in a layer-to-layer connection form [23].
II-B KAN variants
The flexibility of KAN structure allows various implementations using different activation functions on edges. The proposed TaylorKAN and other involved variants are described.
II-B1 TaylorKAN
Taylor series (Eq. 4) can represent a function $f(x)$ as an infinite sum of terms from the values of its derivatives $f^{(n)}$ at a point $a$. TaylorKAN is based on Taylor series, and the coefficients are learned during model training.
$f(x)\approx\sum_{n=0}^{\infty}\frac{f^{(n)}(a)}{n!}(x-a)^{n}$ (4)
II-B2 BSRBF-KAN
BSRBF-KAN combines B-splines (BSs) and radial basis functions (RBFs) to enhance approximation capabilities [26]. BSs ensure the continuity, and RBF provides interpolation. BSRBF can be represented as,
$\phi(x)=w_{b}b(x)+w_{s}(\phi_{BS}(x)+\phi_{RBF}(x)),$ (5)
where $w_{b}$ and $w_{s}$ are the weights of BS and RBF respectively, $\phi_{BS}(x)$ stands for the BS function, and $\phi_{RBF}(x)$ denotes the RBF function. $\phi_{RBF}(r)=e^{-\epsilon r^{2}}$ is typically chosen in which $r$ is the Euclidean distance between the input and the center vector, and $\epsilon$ controls the width of the Gaussian function.
II-B3 ChebyKAN
ChebyKAN employs Chebyshev polynomials for function approximation [27]. It attempts to minimize the maximum error in polynomial approximation and enables ChebyKAN suited for high accuracy and stability. Eq. 6 shows Chebyshev polynomials defined by the recurrence relation.
[…]
II-B4 HermiteKAN
Hermite polynomials are ideal for approximating these Gaussian-like functions due to the recurrence relations and orthogonality. HermiteKAN is based on Hermite polynomials (Eq. 7).
$H_{n}(x)=(-1)^{n}e^{x^{2}}\frac{d^{n}}{dx^{n}}e^{-x^{2}}$ (7)
II-B5 JacobiKAN
JacobiKAN uses Jacobi polynomials. The polynomials are orthogonal to the weight function $(1-x)^{\alpha}(1+x)^{\beta}$ and flexible in handling diverse boundary conditions.
$P_{n}^{(\alpha,\beta)}(x)=\frac{1}{2^{n}}\sum_{k=0}^{n}\binom{n+\alpha}{k}\binom{n+\beta}{n-k}(x-1)^{n-k}(x+1)^{k}$ (8)
II-B6 WavKAN
Wavelets are adept at capturing local variations in functions and provide flexibility and adaptability in complex pattern modeling. The general wavelet transformation form is shown as below,
$\psi_{a,b}(x)=\frac{1}{\sqrt{a}}\psi\left(\frac{x-b}{a}\right),$ (9)
where $a$ and $b$ correspond to the scaling and translation parameters. Since different wavelet basis are available, in this study, the Mexican Hat wavelets are used in WavKAN,
$\psi(x)=\frac{2}{\sqrt{3\pi^{1/4}}}\left(1-x^{2}\right)e^{-\frac{x^{2}}{2}}.$ (10)
II-B7 Other variants
Other KANs are available [24]. However, limited by computing resource, above-mentioned light-weight KANs are evaluated.
III Materials and Methods
III-A Databases
Four databases (BID2011 [30], CID2013 [31], CLIVE [32] and KonIQ-10k [33]) with realistic distortions are analyzed. Table I shows the number (#) of distorted images, the year of data availability, and the score ranges of the databases.
TABLE I: General information of the databases
# images year score range
BID2011 586 2011 [0, 5]
CID2013 474 2013 [0, 100]
CLIVE 1,169 2015 [0, 100]
KonIQ-10k 10,073 2018 [0, 5]
III-B Feature preparation
Two sets of features are prepared. One contains 15 mid-level features per image that are the output of BISA indicators [10]. The other set includes 2048 deeply learned features which are derived from the last full-connection layer of the pre-trained ResNet50 [34]. The procedure is similar as [21].
III-B1 Mid-level features
Assuming a database $\{(I_{i},y_{i})\}_{i=1}^{n}$ with $n$ pairs of images and scores, an indicator $\eta_{j}$ yields an objective score $x_{i,j}$ to an image $I_{i}$, which can be formulated as $x_{i,j}=\eta_{j}(I_{i})$. In the same way, $m$ indicators ($\{\eta_{j}\}_{j=1}^{m}$) generate a feature matrix $M$ as shown in Eq. 11.
$M=\left[\begin{array}[]{ccc|c}x_{1,1}&\dots&x_{1,m}&y_{1}\\ \vdots&\ddots&\vdots&\vdots\\ x_{n,1}&\dots&x_{n,m}&y_{n}\end{array}\right]_{n\times(m+1)}$ (11)
III-B2 High-level features
Using pre-trained ResNet50 as an extractor $f$, it generates a $d$-dimensional feature vector ($d=2048$) to an input image $I_{i}$. Thereby, deeply learned high-level feature matrix $N$ is derived as shown in Eq. 12
$N=\left[\begin{array}[]{c|c}f(I_{1})&y_{1}\\ \vdots&\vdots\\ f(I_{i})&y_{i}\\ \vdots&\vdots\\ f(I_{n})&y_{n}\end{array}\right]_{n\times(d+1)}$ (12)
III-C Score prediction
Besides KANs, SVR and MLP are tested. Assume features ($X=\{\vec{x_{i}}\}_{i=1}^{l}$) and scores ($Y=\{y_{i}\}_{i=1}^{l}$) of $l$ samples in the training set, a weighting vector $\boldsymbol{w}$, and a new input vector $\vec{x}$. SVR ($svr$) aims to find a function $R_{svr}(X)$ that maximizes the deviation of $\epsilon$ from the subjective score $y_{i}$ of training samples. In Eq. 13, $\zeta(X)$ is a nonlinear function, and $\gamma$ is a bias.
$g(X)=R_{svr}(X)=\boldsymbol{w}^{T}\zeta(X)+\gamma$ (13)
MLP ($mlp$) is with one hidden layer of 10 nodes. Its parameters are optimized by minimizing the difference between the model $R_{mlp}$ output and the ground truth $Y$.
$\boldsymbol{w}^{*}=\text{arg min}\quad||Y-R_{mlp}(\boldsymbol{w};X)||^{2}$ (14)
III-D Performance criterion
The performance is evaluated using Pearson linear correlation coefficient (PLCC), Spearman rank order correlation coefficient (SRCC) and Kendalls rank correlation coefficient (KRCC). Higher values indicate better performance.
Notably, PLCC is computed after a five-parameter nonlinear mapping between objective and subjective scores. The mapping function is shown as below,
$f(s)=q_{1}(\frac{1}{2}-\frac{1}{1+e^{q_{2}(s-q_{3})}})+q_{4}s+q_{5},$ (15)
where $s$ denotes the predicted score, $f(s)$ is the mapped score, and $\{q_{i}\}_{i=1}^{5}$ are the fitting parameters. The PLCC value is then calculated between subjective scores and mapped scores.
III-E Implementation details
A database is randomly splitted into three subsets for model training (70% samples), validation (15% samples) and testing (15% samples). All the score prediction models are run on the same data sets.
TaylorKAN is embedded with 2 hidden layers ([15, 8, 4, 1]) for mid-level feature input and with 3 hidden layers ([2048, 512, 256, 128, 1]) for high-level feature input. Specifically, quadratic approximation is used in TaylorKAN. Other models are used with default parameter settings. The project code is available online (www.github.com/CUC-Chen/kanBISA).
IV Results and Discussion
Score prediction results using mid-level features and high-level features are shown in Table II and Table III, respectively. Table IV compares the current study with several state-of-the-art achievement. The highest metric values are in boldface.
IV-A BISA using mid-level features
Table II shows that compared to SVR, all the KAN models are better on BID2011, TaylorKAN and HermiteKAN are superior on CID2013, TaylorKAN, BSRBF_KAN, JacobiKAN, and HermiteKAN achieve slightly higher PLCC on CLIVE, and the KANs lead to competitive results on KonIQ-10k.
TABLE II: No-reference assessment results by using mil-level features
BID2011 CID2013 CLIVE KonIQ-10k
PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC
SVR 0.564 0.561 0.391 0.831 0.822 0.683 0.589 0.561 0.392 0.764 0.732 0.535
MLP 0.589 0.588 0.418 0.839 0.835 0.681 0.572 0.542 0.381 0.752 0.721 0.526
BSRBF_KAN 0.697 0.705 0.519 0.833 0.802 0.600 0.601 0.514 0.360 0.742 0.667 0.481
ChebyKAN 0.711 0.769 0.560 0.753 0.792 0.606 0.522 0.454 0.306 0.752 0.675 0.488
HermiteKAN 0.721 0.779 0.579 0.815 0.822 0.634 0.581 0.522 0.368 0.747 0.678 0.490
JacobiKAN 0.719 0.778 0.567 0.792 0.830 0.631 0.572 0.475 0.324 0.749 0.676 0.489
WavKAN 0.655 0.685 0.486 0.850 0.824 0.641 0.606 0.498 0.343 0.762 0.690 0.502
TaylorKAN 0.757 0.750 0.550 0.859 0.843 0.666 0.608 0.488 0.342 0.719 0.640 0.458
It is observed that when using 15 mid-level features, the proposed TaylorKAN leads to the highest PLCC on BID2011, CID2013 and CLIVE, and slightly lower values on KonIQ-10k to the baseline. The reasons are manifold. Firstly, mid-level features intrinsically represent image sharpness. And thereby, all KANs, particularly TaylorKAN, obtain much higher values on BID2011. Secondly, the characteristic of image sharpness representation constrains the prediction performance of KANs on CID2013, CLIVE and KonIQ-10k, since these databases contains diverse distortions. It becomes understandable that KANs obtain inferior results on the largest, realistic database KonIQ-10k. Thirdly, the representation capacity of the mid-level features might be insufficient that poses difficulties on accurate score prediction.
IV-B BISA using deeply learned features
Setting SVR as the baseline model, Table III shows KANs achieve higher values on BID2011 and much lower values on CID2013, BSRBF_KAN and WavKAN obtain slight increase on PLCC on CLIVE, and KANs achieve superior performance on KonIQ-10k.
TABLE III: No-reference assessment results by using deeply learned features
BID2011 CID2013 CLIVE KonIQ-10k
PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC
SVR 0.774 0.752 0.562 0.924 0.909 0.745 0.733 0.724 0.533 0.803 0.783 0.585
MLP 0.746 0.725 0.538 0.925 0.909 0.748 0.465 0.440 0.310 0.607 0.576 0.408
BSRBF_KAN 0.818 0.813 0.607 0.858 0.861 0.689 0.737 0.663 0.487 0.840 0.807 0.615
ChebyKAN 0.809 0.816 0.621 0.525 0.527 0.382 0.683 0.545 0.394 0.838 0.804 0.611
HermiteKAN 0.825 0.831 0.635 0.636 0.635 0.472 0.741 0.704 0.520 0.844 0.810 0.617
JacobiKAN 0.826 0.834 0.633 0.683 0.685 0.513 0.747 0.694 0.515 0.842 0.809 0.616
WavKAN 0.797 0.811 0.608 0.839 0.850 0.663 0.761 0.687 0.507 0.848 0.817 0.625
TaylorKAN 0.792 0.788 0.587 0.750 0.756 0.565 0.667 0.563 0.405 0.831 0.801 0.610
When using 2048 high-level features for image quality representation, TaylorKAN obtains slightly higher metric values on BID2011 and KonIQ-10k, and lower values on CID2013 and CLIVE to SVR. The inferior performance suggests that TaylorKAN may not be well-suited for handling high-dimensional features. Compared to the metric values derived from mid-level features, all KANs, except BSRBF_KAN, show significant decrease on CID2013, and their improvements on the other databases are not as substantial as those of SVR. Another reason might come from the representation capacity of quadratic approximation in TaylorKAN. Higher-order Taylor expansions may improve prediction accuracy, while simpler approximations can simplify complex problems, enhance the efficiency of the optimization process, and reduce the risk of over-fitting in score prediction.
IV-C Current achievement on the databases
The results of TaylorKANs using mid-level features (a) and high-level features (b) and several state-of-the-art works are shown in Table IV. These works [15, 14, 16, 17, 18, 19, 20] design novel CNNs and take images as input rather than features.
TABLE IV: State-of-the-art performance on realistic distorted image datasets
BID2011 CID2013 CLIVE KonIQ-10k
PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC
TaylorKAN a 0.757 0.750 0.550 0.859 0.843 0.666 0.608 0.488 0.342 0.719 0.640 0.458
TaylorKAN b 0.792 0.788 0.587 0.750 0.756 0.565 0.667 0.563 0.405 0.831 0.801 0.610
[14] 0.861 0.846 0.934 0.930 0.873 0.855 0.923 0.901
[15] 0.873 0.858 0.890 0.854 0.901 0.896
[16] 0.886 0.874 0.880 0.865 0.916 0.904
[17] 0.891 0.875 0.898 0.882 0.921 0.912
[18] 0.9284 0.9128 0.9175 0.8992 0.9362 0.9209
[19] 0.9109 0.8875 0.9141 0.8947 0.9413 0.9271
[20] 0.900 0.875 0.910 0.904 0.908 0.919
A significant performance gap exists between CNNs and TaylorKAN in score prediction. CNNs achieve PLCC $\geq$ 0.86 and SRCC $\geq$ 0.84, which are notably higher than those of the proposed KAN. These works benefit not only from advanced architectures of image convolution and feature pooling but also from sophisticated module designs, such as hinge constraints on learning uncertainty [15], multi-scale spatial pyramid pooling [17], self-supervised learning [19], and multitask learning [20]. Therefore, it becomes urgent to integrate prior knowledge and advanced modules into KANs for improving representation power and score prediction performance [23, 24].
When using TaylorKAN for score prediction, deeply learned features can increase, but not substantially, metric values on three databases. On the one hand, since perceptual similarity is an emergent property shared across deep visual representation [35], it is not surprise that deep features are effective for image quality representation [21, 19]. On the other hand, selecting the most informative subsets of deep features is urgently important for KANs to enhance accuracy and efficiency in score prediction, while avoiding the challenges of high-dimensional feature processing.
V Conclusions
Besides SVR and MLP, six KANs are evaluated on four realistic databases for score prediction using mid-level features and high-level features. KANs achieve superior or competitive results over SVR and MLP, showing promise on improving the performance on score prediction and related tasks.